People commonly dissolve table salt (NaCL) into water that they are heating up.

a) If you had a pot with 4 quarts of water (before heating) and add 1 teaspoon of salt, which is equivalent to 6.20g, what is the boiling point of the water? How much of an effect will that have on cooking pasta? (You can assume that NaCl fully dissociates. 1quart=0.946 L and Kb=0.51oCkg/mol for water)

b) Would the cooking of pasta (based on boiling point) be affected more or less if one added the same concentration of table sugar (sucrose C12H22O11)? Explain
Answer #1

a ) DTb = i*Kb*m

i = 2 for NaCl
m = molality of solution = w/mwt*1/wt of solvent

          = (6.2/58.5)*(1/(4*0.946))         (water density = 1 g/cc. volume = weight)

      = 0.028 molal

Kb of water = 0.51oCkg/mol

Boilining point of water = 100 C

Tb -100 = 2* 0.51*(0.028)

Tb = boiling point of solution = 100.0286 C

b) if the table sugar is used

i = 1 sugar

Tb -100 = 1*0.51*(0.028)

Tb = 100.0143 C

so that by increasing boiling point of solution the cooking of pasta become more easy.

there for usage of salt make more easier than sugar.